OpenStudy (keke):

find the vertex of the quadratic function: f(x)=3x^(2)-12x=16

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

2,4

OpenStudy (anonymous):

to find the vertex either "complete the square" or just use \[x=-\frac{b}{2a}\] in this case b = -12, a = 3 so \[-\frac{b}{2a}=-\frac{-12}{2\times 3} = 2\] to get the y value find \[f(2)=3\times 2^2-12\times 2 -16=12-24-16=-28\]

OpenStudy (anonymous):

btw i am assuming this is \[f(x)=3x^2-12x-16\] not \[f(x)=3x^2-12x=16\] which doesn't make any sense